# Phi-4-Mini Technical Report: Compact yet Powerful Multimodal Language Models via Mixture-of-LoRAs

### v1

@@ 2.2.2 Training Pipeline
Recent studies have suggested that training a robust reasoning model only requires a small amount of high-quality data, such as LIMO [YHX+25] and S1K [MYS+25]. However, we propose a fundamentally different training paradigm for SLM: first, we need to conduct a pre-training phase on extensive reasoning data to capture general reasoning chains, and then perform careful fine-tuning on curated SFT or preference data. The continued training of Phi-4-Mini for reasoning proceeds in three distinct stages. First, building on Phi-4-Mini, the model is pre-trained on approximately 60 billion reasoning CoT tokens generated by frontier reasoning LLMs, after which rejection sampling is employed to filter out incorrect outputs. This allows the reasoning extension of Phi-4-Mini to learn the reasoning chains produced by these models. In the second stage, the model is fine-tuned on a smaller but carefully curated dataset of around 200K high-quality CoT samples, chosen to cover diverse domains and varying difficulty levels. Finally, in the third stage, we label filtered incorrect outputs as “dis-preferred” and their corrected counterparts as ‘preferred’, compiling a new dataset of 300K preference samples for DPO training.
@@ 4.2.2 Coding
Model AIME MATH-500 GPQA Diamond o1-mini* 63.6 90.0 60.0 DeepSeek-R1-Distill-Qwen-7B 53.3 91.4 49.5 DeepSeek-R1-Distill-Llama-8B 43.3 86.9 47.3 Bespoke-Stratos-7B* 20.0 82.0 37.8 OpenThinker-7B* 31.3 83.0 42.4 Llama-3.2-3B-Instruct 6.7 44.4 25.3 Phi-4-Mini 10.0 71.8 36.9 Phi-4-Mini (reasoning trained) (3.8B) 50.0 90.4 49.0
@@ 4.2.3 CoT Reasoning
We evaluate the reasoning performance of a reasoning-enhanced model that we have trained over Phi-4-Mini. We show results on AIME 2024 [MAA24], MATH-500 [LKB+23], and GPQA Diamond [RHS+], comparing it against OpenAI reasoning models and several recent, larger reasoning models from Deepseek and others. Despite having only 3.8B parameters, Phi-4-Mini reasoning-enhanced model outperforms DeepSeek-R1-Distill-Llama-8B [GYZ+25], Bespoke-Stratos-7B [Lab25], OpenThinker-7B [Tea25a], and achieves performance comparable to DeepSeek-R1-Distill-Qwen-7B as shown in the Table 9.
@@ Appendix B Authors (alphabetical)
Xiyang Dai Ali Mousavi Ziyi Yang
Ruchao Fan Anh Nguyen Donghan Yu
Mei Gao Jing Pan Ishmam Zabir
Min Gao Daniel Perez-Becker Jianwen Zhang
Amit Garg Jacob Platin Li Lyna Zhang
Abhishek Goswami Thomas Portet Yunan Zhang
Junheng Hao Kai Qiu Xiren Zhou
Amr Hendy

### v2

@@ 2.2.2 Training Pipeline
Recent studies have suggested that training a robust reasoning model only requires a small amount of high-quality data, such as LIMO [YHX+25] and S1K [MYS+25]. However, we propose a fundamentally different training paradigm for SLM: we need to conduct a pre-training phase on extensive reasoning data to capture general reasoning chains, and then perform careful fine-tuning on curated SFT or preference data. The continued training of Phi-4-Mini for reasoning proceeds in three distinct stages. 1) First, building on Phi-4-Mini, the model is pre-trained on approximately 60 billion reasoning CoT tokens generated by frontier reasoning LLMs, after which rejection sampling is employed to filter out incorrect outputs. This allows the reasoning extension of Phi-4-Mini to learn the reasoning chains produced by these models. 2) In the second stage, the model is fine-tuned on a smaller but carefully curated dataset of around 200K high-quality CoT samples, chosen to cover diverse domains and varying difficulty levels. 3) Roll-Out DPO: finally, in the third stage, we label filtered incorrect outputs as “dis-preferred” and their corrected counterparts as ‘preferred’, compiling a new dataset of 300K preference samples for DPO training.
@@ 4.2.2 Coding
Model AIME MATH-500 GPQA Diamond o1-mini* 63.6 90.0 60.0 DeepSeek-R1-Distill-Qwen-7B 53.3 91.4 49.5 DeepSeek-R1-Distill-Llama-8B 43.3 86.9 47.3 Bespoke-Stratos-7B* 20.0 82.0 37.8 OpenThinker-7B* 31.3 83.0 42.4 Llama-3.2-3B-Instruct 6.7 44.4 25.3 Phi-4-Mini 10.0 71.8 36.9    + Distill Pre-training 30.0 82.9 42.6    + Distill Fine-tuning 43.3 89.3 48.3    + Roll-Out DPO (final reasoning-enhanced Phi-4-Mini) 50.0 90.4 49.0
@@ 4.2.3 CoT Reasoning
We evaluate the reasoning performance of a reasoning-enhanced model that we have trained over Phi-4-Mini. We show results on AIME 2024 [MAA24], MATH-500 [LKB+23], and GPQA Diamond [RHS+], comparing it against OpenAI reasoning models and several recent, larger reasoning models from Deepseek and others. Despite having only 3.8B parameters, Phi-4-Mini reasoning-enhanced model outperforms DeepSeek-R1-Distill-Llama-8B [GYZ+25], Bespoke-Stratos-7B [Lab25], OpenThinker-7B [Tea25a], and achieves performance comparable to DeepSeek-R1-Distill-Qwen-7B as shown in the Table 9. Moreover, we present an ablation study that shows the effectiveness of our training process for reasoning-enhanced Phi-4-Mini.
@@ Appendix B Authors (alphabetical)
Xiyang Dai Arindam Mitra Ziyi Yang
Ruchao Fan Ali Mousavi Donghan Yu
Mei Gao Anh Nguyen Ishmam Zabir
Min Gao Jing Pan Jianwen Zhang
Amit Garg Daniel Perez-Becker Li Lyna Zhang
Abhishek Goswami Jacob Platin Yunan Zhang
Junheng Hao Thomas Portet Xiren Zhou
Amr Hendy Kai Qiu